OpenStudy (anonymous):

What is the value of a in the equation a = 2 + 3a + 8?

pooja195 (pooja195):

a=2+3a+8 a=(3a)+(2+8)(Combine Like Terms) a=3a+10 a=3a+10 Subtract 3a from both sides. a−3a=3a+10−3a −2a=10 Finish it off.

OpenStudy (anonymous):

Where did you get 3a from?

OpenStudy (anonymous):

i mean 2a

pooja195 (pooja195):

a-3a=-2a

OpenStudy (anonymous):

im so dumb sorry

pooja195 (pooja195):

lol no you are not :)

pooja195 (pooja195):

Alright so finish it off! :D

OpenStudy (anonymous):

sorry I'm so slow! I'm not good at math

pooja195 (pooja195):

lol its alright −2a=10 divide both sides by -2 :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

a=-5

pooja195 (pooja195):

Yep :D

OpenStudy (anonymous):

thank you so much!

pooja195 (pooja195):

No problem! :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

:)